1-12 Martinez Company’s relevant range of production is 7,500 units to 12,500 units. When it produces...

1-12

Martinez Company’s relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows:

Average Cost per Unit
Direct materials $ 5.20
Direct labor $ 2.70
Variable manufacturing overhead $ 1.50
Fixed manufacturing overhead $ 4.00
Fixed selling expense $ 2.20
Fixed administrative expense $ 2.00
Sales commissions $ 1.00
Variable administrative expense $ 0.50

If 12,500 units are produced, what is the total amount of manufacturing overhead cost incurred to support this level of production? What is this total amount expressed on a per unit basis? (Round your "per unit" answer to 2 decimal places and other answers to the nearest whole dollar amount.)
Answer #1

Manufacturing Overhead cost per unit

= Total Variable + Total fixed

= (12,500*1.50) + (10,000*4) = 58,750

Per unit = 58,750/12,500

= 4.70